OpenStudy (anonymous):

How much work would be required to move an electron from 0.0529 × 10-9 m from a proton to four times that distance? 5.4 × 10-18J 3.3 × 10-18J 1.6 × 10-18J None are correct.

OpenStudy (experimentx):

k q^2/(0.0529 × 10-9 ) * (1 - 1/4)

OpenStudy (anonymous):

I don't understand can you write the rule

OpenStudy (mos1635):

E1=E2 (mechanikal energy) K1+U1=K2+U2 U1-U2=K2-K1 U1-U2=ΔΚ U1-U2=W so W=k*Q1*Q2/R1-k*Q1*Q2/R2

OpenStudy (mos1635):

Q1=Q2=q R2=4*R1 so W=k*q^2/R*(1-1/4) .......